The length of a rectangle is 4 cm more than 2 times its width. If the area of the rectangle is 74 cm^2, find the dimensions of the rectangle to the nearest thousandth.
Answer by checkley71(8403)

L=2W+4
W(2W+4)=74
2W^2+4W-74=0
2(W^2+2W-37)=0
USING THE QUADRATIC EQUATION:
W=(-2+0-SQRT[2^2-4*1*-37])/2*1
W=(-2+-SQRT[4+148])/2
W=(-2+-SQRT[152)/2
W=(-2+-12.3288)/2
W=(-2+12.3288)/2
W=10.3288/2
W=5.164 CM. ANSWER FOR THE WIDTH.
L=2*5.164+4
L=10.328+4
L=14.328 ANSWER FOR THE LENGTH.
PROOF
5.164*14.328=74
74~74